$SPAD/src/input bug9057.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool bug9057.output
)set message test on
)set message auto off
)clear all
--S 1 of 9
g:=operator 'g
--R
--R
--R   (1)  g
--R                                                          Type: BasicOperator
--E 1
--S 2 of 9
f==n+->sum(g(j),j=1..n)
--R
--R                                                                   Type: Void
--E 2
--S 3 of 9
f(1)
--R
--R   Compiling function f with type PositiveInteger -> Expression(Integer
--R      )
--R
--R   (3)  g(1)
--R                                                    Type: Expression(Integer)
--E 3
--S 4 of 9
f==n+->product(sum(1/i,i=1..j),j=1..n)
--R
--R   Compiled code for f has been cleared.
--R   1 old definition(s) deleted for function or rule f
--R                                                                   Type: Void
--E 4
--S 5 of 9
f(1)

--R
--R   Compiling function f with type PositiveInteger -> Expression(Integer
--R      )
--R
--R   (5)  1
--R                                                    Type: Expression(Integer)
--E 5
--S 6 of 9
f==n+->product(product(1/i,i=1..j),j=1..n)
--R
--R   Compiled code for f has been cleared.
--R   1 old definition(s) deleted for function or rule f
--R                                                                   Type: Void
--E 6
--S 7 of 9
f(1)
--R
--R   Compiling function f with type PositiveInteger -> Expression(Integer
--R      )
--R
--R   (7)  1
--R                                                    Type: Expression(Integer)
--E 7
--S 8 of 9
f==n+->sum(sum(1/i,i=1..j),j=1..n)
--R
--R   Compiled code for f has been cleared.
--R   1 old definition(s) deleted for function or rule f
--R                                                                   Type: Void
--E 8
--S 9 of 9
f(1)
--R
--R   There are 6 exposed and 2 unexposed library operations named sum
--R      having 2 argument(s) but none was determined to be applicable.
--R      Use HyperDoc Browse, or issue
--R                               )display op sum
--R      to learn more about the available operations. Perhaps
--R      package-calling the operation or using coercions on the arguments
--R      will allow you to apply the operation.
--R   Cannot find a definition or applicable library operation named sum
--R      with argument type(s)
--R          Union(Fraction(Polynomial(Integer)),Expression(Integer))
--R                       SegmentBinding(PositiveInteger)
--R
--R      Perhaps you should use "@" to indicate the required return type,

--R      or "$" to specify which version of the function you need.
--R   AXIOM will attempt to step through and interpret the code.
--R
--R   (9)  1
--R                                                    Type: Expression(Integer)
--E 9
)spool
)lisp (bye)

———-

References^{[1] nothing}

6